Please help me solve this rational equation:

x/(x-1)+ x = 4x-3/x-1

Thanks

Answer by oscargut(686)

x/(x-1)+ x = 4x-3/x-1 if you multiply by (x-1) you have:
x+x(x-1)=4x-3
x^2-4x+3=0 solutions are 1 and 3 but 1 is not a solution of the firs equation
so solution is x=3
Answer: x=3